Part A. What would be the luminosity of the Sun if its surface temperature were 3900 K and its radius were 1.5 AU ?

Part B. What would be the luminosity of the Sun if its surface temperature were 3900 K and its radius were 7.0 AU ?

*I don't understand how to calculate this out. Instructor didn't really cover this.
Answer #1

When speaking of the Luminosity, it can best be described as the total amount of light that will be emitted from the Surface of the sun, in a given period of time. the Luminotsity of the star, or the Sun (as in this question) can be calculated using different equations (in genral termed as the Luminosity Equation).

if we have to calculate the Luminosity odf the interstellar star,then it is calcualetd on the basis of comparison between the Luminosity of the Sun (Radius and surface temperature ratio).

but here in thiscase, we arer only to calculate the Luminosity of the Sun's surface, so we consider using simple formula.

A) First we have - Radius = 1.5 AU, and the ST (Surface Temp.) = 3900 K.

so, the L = R2T4

L = (1.5)2(3900)4 = 2.25 x (39 x 102)4 = 2.25 x (23,13,441 x 108)= 52,05,242 x 10^8. K.A.U.

B) Second we have - Radius = 7 AU, and Temp = 3900 K

usig the same formula we have (7)2 x (39 x 102)4 = 49 x 23,13,441 x 108 = 1,13,35 x 108 K. AU.
Answer #2

Part A: Luminosity at 3900 K and 1.5 AU

  1. Convert radius to meters:

    R=1.5AU=1.5×1.496×1011m=2.244×1011m

  2. Plug into the formula:

    L=4π(2.244×1011)2×(5.67×108)×(3900)4

  3. Simplify step-by-step:

    • Calculate R2:

      (2.244×1011)2=5.036×1022m2

    • Calculate T4:

      (3900)4=2.31×1014K4

    • Multiply all terms:

      L=4π×5.036×1022×5.67×108×2.31×1014L8.23×1030W

  4. Compare to the Sun's current luminosity:

    LL=8.23×10303.828×102621,500

  5. Answer: The Sun would be ~21,500 times more luminous than its current state.



Part B: Luminosity at 3900 K and 7.0 AU

  1. Convert radius to meters:

    R=7.0AU=7.0×1.496×1011=1.047×1012m

  2. Plug into the formula:

    L=4π(1.047×1012)2×(5.67×108)×(3900)4

  3. Simplify step-by-step:

    • R2:

      (1.047×1012)2=1.096×1024m2

    • T4: Same as Part A (2.31×1014K4).

    • Multiply all terms:

      L=4π×1.096×1024×5.67×108×2.31×1014L1.79×1032W

  4. Compare to the Sun's current luminosity:

    LL=1.79×10323.828×1026467,000

  5. Answer: The Sun would be ~467,000 times more luminous than its current state.



answered by: anonymous